# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ Abstract.
In this paper, we study two types of inverse problems for space semidiscrete stochastic parabolic equations in arbitrary dimensions. The first problem concerns a semidiscrete inverse source problem, which involves determining the random source term of the white noise in the stochastic parabolic equation using observation data of the solution at the terminal time and the trace of the discrete spatial derivative of the solution on a subdomain at the boundary of the space. The second problem addresses a semidiscrete Cauchy inverse problem, which involves determining the solution of the stochastic parabolic equation in a subdomain of space-time, from the data of the solution and the trace of the discrete spatial derivative on a subdomain of the boundary of space and time. To address these problems, we establish two Carleman estimates for a semidiscrete stochastic parabolic operator in arbitrary dimensions: one for the homogeneous boundary condition and the other for the nonhomogeneous boundary condition. Applying these Carleman estimates, we obtain Lipschitz and Hölder stability for the first and second inverse problems, respectively.
@@
semidiscrete stochastic parabolic equations, global Carleman estimate, Inverse problems.
@@ 1. Introduction
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a complete filtered probability space on which a one-dimensional standard Brownian motion $\{B(t)\}_{t\geq 0}$ is defined such that $\{\mathcal{F}_{t}\}_{t\geq 0}$ is the natural filtration generated by $B(t)$, augmented by all $\mathbb{P}$ null sets in $\mathcal{F}$ and $\mathcal{X}$ a Banach space. The space $L^{2}_{\mathcal{F}}(\Omega;L^{2}(0,T;\mathcal{X}))$ consists of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted processes $X(\cdot)$ such that $\mathbb{E}\left(\int_{0}^{T}\|X(\cdot)\|^{2}_{\mathcal{X}}\,dt\right)<\infty.$ The space $L^{2}_{\mathcal{F}}(\Omega;L^{\infty}(0,T;\mathcal{X}))$ consists of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted bounded processes. Finally, $L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathcal{X}))$ denotes the space of all $L^{p}_{h}(\mathcal{M})$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted continuous processes $X(\cdot)$ such that $\mathbb{E}\left(\sup_{t\in[0,T]}\|X(\cdot)\|^{2}_{\mathcal{X}}\right)<\infty.$ For simplicity, we often write $L^{2}_{\mathcal{F}}(0,T;\mathcal{X})$ or $L^{\infty}_{\mathcal{F}}(0,T;\mathcal{X})$ whenever no confusion may arise.
Let $T>0$ and $n\in\mathds{N}$, and $G\triangleq(0,1)^{n}$. We consider the following stochastic parabolic equation:
@@
In (1.1), $y$ denotes the state variable, $g$ is a source term representing the intensity of the random force of white noise, and $\gamma^{i},a^{1i}$, and $a^{k}$ (for $i=1,\ldots,n$ and $k=2,3$) are given functions. We refer to KR77 , Zho92 , and lu2021mathematical for the existence and uniqueness results of (1.1).
The main objective of this paper is to study the inverse random source problem and the Cauchy problem associated with the semidiscrete spatial approximation of the stochastic parabolic equation (1.1). To this end, we introduce the spatial discretization employed throughout this work.
Given a set of points $\mathcal{M}\subseteq G\cap h\mathbb{Z}^{n}$, called the primal mesh, we define the dual mesh in the direction $e_{k}$, where $\{e_{k}\}_{k=1}^{n}$ denotes the standard basis of $\mathbb{R}^{n}$, by
$\mathcal{M}_{k}^{\ast}\triangleq\left\{x\pm\frac{h}{2}e_{k}\mid x\in\mathcal{M}\right\}.$
This allows us to define the boundary points in the $e_{k}$ direction as $\partial_{k}\mathcal{M}\triangleq(\mathcal{M}^{\ast}_{k})^{\ast}_{k}\setminus\mathcal{M}$, and we set
$\partial\mathcal{M}\triangleq\bigcup_{k=1}^{n}\partial_{k}\mathcal{M},\quad\textrm{ and }\quad\overline{\mathcal{M}}\triangleq\bigcup_{k=1}^{n}(\mathcal{M}^{\ast}_{k})^{\ast}_{k}.$
@@
(1.2) $\forall x\in\partial_{i}\mathcal{M},\quad\nu_{i}(x)\triangleq\begin{cases}1&\mbox{if }x-\tfrac{h}{2}e_{i}\in\mathcal{M}_{i}^{\ast}\ \mbox{and }x+\tfrac{h}{2}e_{i}\notin\mathcal{M}_{i}^{\ast},\\ -1&\mbox{if }x-\tfrac{h}{2}e_{i}\notin\mathcal{M}_{i}^{\ast}\ \mbox{and }x+\tfrac{h}{2}e_{i}\in\mathcal{M}_{i}^{\ast},\\ 0&\mbox{otherwise}.\end{cases}$
@@
(1.3) $\forall x\in\partial_{i}\mathcal{M},\quad t_{r}^{i}(u)(x)\triangleq\begin{cases}u\!\left(x-\tfrac{h}{2}e_{i}\right)&\nu_{i}(x)=1,\\ u\!\left(x+\tfrac{h}{2}e_{i}\right)&\nu_{i}(x)=-1,\\ 0&\nu_{i}(x)=0.\end{cases}$
@@
$(D_{k}u)(x)\triangleq\frac{\tau_{+k}u(x)-\tau_{-k}u(x)}{h},\quad(A_{k}u)(x)\triangleq\frac{\tau_{+k}u(x)+\tau_{-k}u(x)}{2},\quad\text{for }x\in\mathcal{M}^{\ast}_{k},$
where $\tau_{\pm k}u(x)\triangleq u\!\left(x\pm\frac{h}{2}e_{k}\right)$.
@@
$\|u\|_{L_{h}^{p}(\mathcal{M})}^{p}\triangleq\int_{\mathcal{M}}|u|^{p},$
where the integral is defined by
$\int_{\mathcal{M}}|u|^{p}\triangleq h^{n}\sum_{x\in\mathcal{M}}|u(x)|^{p}.$
Analogously, we denote by $W^{1,p}_{h}(\mathcal{M})$, $W_{h}^{2,p}(\mathcal{M})$, and $L^{\infty}_{h}(\mathcal{M})$ the set of functions defined in $\mathcal{M}$ with the associated norms.
$\|u\|_{W_{h}^{1,p}(\mathcal{M})}^{p}\triangleq\|u\|_{L_{h}^{p}(\mathcal{M})}^{p}+\sum_{i\in\llbracket 1,n\rrbracket}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}u|^{p},$
$\|u\|_{W_{h}^{2,p}(\mathcal{M})}^{p}\triangleq\|u\|_{W_{h}^{1,p}(\mathcal{M})}^{p}+\sum_{i,j\in\llbracket 1,n\rrbracket}\int_{Q_{ij}^{\ast}}|D_{ij}^{2}u|^{p},$
and
$\|u\|_{L_{h}^{\infty}(\mathcal{M})}\triangleq\max_{x\in\mathcal{M}}|u(x)|,$
@@
$\int_{\partial\mathcal{M}}u\triangleq\sum_{i=1}^{n}\int_{\partial_{i}\mathcal{M}}u(x),$
where
$\int_{\partial_{i}\mathcal{M}}u(x)\triangleq h^{n-1}\sum_{x\in\partial_{i}\mathcal{M}}u(x).$
Then, we can denote as $L^{2}_{h}(\partial\mathcal{M})$ and $H^{1}_{h}(\partial\mathcal{M})$ the set of functions defined on $\partial\mathcal{M}$ with the associated norm
$\left\|u\right\|^{2}_{L^{2}_{h}(\partial\mathcal{M})}\triangleq\int_{\partial\mathcal{M}}|u|^{2},$
and
$\left\|u\right\|^{2}_{H^{1}_{h}(\partial\mathcal{M})}\triangleq\sum_{i\in\llbracket 1,n\rrbracket}\left(\int_{\partial_{i}\mathcal{M}}|u|^{2}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\int_{(\partial_{i}\mathcal{M})_{j}^{\ast}}|D_{j}u|^{2}\right),$
respectively. Finally, following ervedoza-2011 , we adopt the discrete $H^{1/2}_{h}$-norm defined for $v\in C(\partial\mathcal{M})$ by
$\|v\|_{H_{h}^{1/2}(\partial\mathcal{M})}\triangleq\min_{\left.u\right|_{\partial\mathcal{M}}=v}\|u\|_{H^{1}_{h}(\mathcal{M})}.$
Following the same reasoning, we analogously define the semidiscrete spaces $L^{2}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\partial\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M}))$, and $L^{2}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))$.
Given the previous definitions, we are now ready to formulate the problems to be addressed. Let $\gamma_{i}$ denote the sampling of $\gamma^{i}$ onto both the primal and dual meshes. Furthermore, let $a_{1i}$, $a_{2}$, $a_{3}$, and $a_{4}$ represent the evaluations of $a^{1i}$, $a^{2}$, and $a^{3}$ on the primal mesh, for $i=1,\ldots,n$, respectively. We now formulate the discrete versions of the inverse problems considered in this paper, namely, the discrete inverse random source problem and the discrete Cauchy problem for the spatially semidiscrete stochastic parabolic equation in $n$ dimensions.
Semidiscrete inverse random source problem. We consider the following $n$-dimensional semidiscrete stochastic parabolic equation, derived from (1.1), assuming $a_{3}=0$:
@@
where the second-order operator is defined by $\mathcal{A}_{h}w\triangleq dw-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(w))dt$, and $w_{0}$ is the initial condition.
We define the observation operator $\Lambda_{1}:L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))\longrightarrow X_{1}$ as
$\Lambda_{1}(g)\triangleq\left(\left.\partial_{\nu}w\right|_{\Gamma^{+}\times(0,T)},w(T)\right),$
where the normal derivative is given by
$\partial_{\nu}w\triangleq\sum_{i=1}^{n}t_{r}^{i}(D_{i}w)e_{i}\nu_{i}.$
The set $\Gamma^{+}\subset\partial\mathcal{M}$ is defined as
$\Gamma^{+}\triangleq\left\{x\in\partial\mathcal{M}:x\in\partial_{i}\mathcal{M}\text{ and }(x-x^{*})e_{i}\nu_{i}\geq 0\right\},$
for some fixed point $x^{*}\notin\overline{\mathcal{M}}$, and the norm that will be adopted in $X_{1}\triangleq\left(L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))\right)^{n}\times L^{2}_{\mathcal{F}_{T}}(\Omega;L_{h}^{2}(\mathcal{M}))$ is
$\left\|\left(f_{1},...,f_{n},f_{n+1}\right)\right\|_{X_{1}}\triangleq\sum_{i=1}^{n}\|f_{i}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}_{h}(\Gamma^{+}))}+\|f_{n+1}\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;L^{2}_{h}(\mathcal{M}))},$
for $f_{i}\in L^{2}_{\mathcal{F}}(0,T;L^{2}_{h}(\mathcal{M}))$ with $i=1\cdots n$ and $f_{n+1}\in L^{2}_{\mathcal{F}_{T}}(\Omega;L^{2}_{h}(\mathcal{M}))$.
The direct problem consists of solving (1.4) for a given $g$ to compute $\Lambda_{1}(g)$. The inverse problem is to recover $g$ from the knowledge of $\Lambda_{1}(g)$.
We present a conditional stability result.
@@
Let $a_{1i}\in L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$, $a_{2}\in L_{\mathcal{F}}^{\infty}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))$ where the exponent $n^{\ast}$ satisfies
@@
and assume that $g_{k}\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))$ for $k=1,2$, which satisfy
(1.6) $|D_{i}(g_{1}-g_{2})|\leq C|A_{i}(g_{1}-g_{2})|,\quad\text{in }Q_{i}^{*},\;\mathbb{P}\text{-a.s.},\quad\text{for }i=1,\dots,n.$
Then, there exists a constant $C>0$, independent of $h$, such that
(1.7) $\|g_{1}-g_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\mathcal{M}))}\leq C\|\Lambda_{1}(g_{1})-\Lambda_{1}(g_{2})\|_{X_{1}}.$

### v2

# Inverse Random Source and Cauchy Problems for Semi-Discrete Stochastic Parabolic Equations in Arbitrary Dimensions

@@ Abstract.
In this paper, we study two types of inverse problems for space semi-discrete stochastic parabolic equations in arbitrary dimensions. The first problem concerns a semi-discrete inverse source problem, which involves determining the random source term of the white noise in the semi-discrete stochastic parabolic equation using observation data of the solution at the terminal time and on an arbitrary open spatial subdomain over a time interval. The second problem addresses a semi-discrete Cauchy inverse problem, which involves determining the solution of the stochastic parabolic equation in a space-time subdomain, from measurements of the solution and the trace of its discrete spatial derivative on an arbitrary open subset of the lateral boundary over a time interval. The key tools are three new global Carleman estimates for the semi-discrete stochastic parabolic operator, one for interior observations and two for boundary observations (homogeneous and nonhomogeneous Dirichlet conditions). Applying these Carleman estimates, we obtain Lipschitz and Hölder stability for the first and second inverse problems, respectively.
@@
Inverse problems, Stochastic processes, semi-discrete stochastic parabolic equations, global Carleman estimate.
@@ 1. Introduction
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a complete filtered probability space on which a one-dimensional standard Brownian motion $\{B(t)\}_{t\geq 0}$ is defined such that $\{\mathcal{F}_{t}\}_{t\geq 0}$ is the natural filtration generated by $B(t)$, augmented by all $\mathbb{P}$ null sets in $\mathcal{F}$ and $\mathcal{X}$ a Banach space. The space $L^{2}_{\mathcal{F}}(\Omega;L^{2}(0,T;\mathcal{X}))$ consists of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted processes $X(\cdot)$ such that $\mathbb{E}\left(\int_{0}^{T}\|X(\cdot)\|^{2}_{\mathcal{X}}\,dt\right)<\infty.$ The space $L^{2}_{\mathcal{F}}(\Omega;H^{1}(0,T;\mathcal{X}))$ consists of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted processes $X(\cdot)$ such that $\mathbb{E}\left(\int_{0}^{T}\|X_{t}(\cdot)\|^{2}_{\mathcal{X}}+\|X(\cdot)\|^{2}_{\mathcal{X}}\,dt\right)<\infty.$ The space $L^{2}_{\mathcal{F}}(\Omega;L^{\infty}(0,T;\mathcal{X}))$ consists of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted bounded processes. Finally, $L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathcal{X}))$ denotes the space of all $\mathcal{X}$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted continuous processes $X(\cdot)$ such that $\mathbb{E}\left(\sup_{t\in[0,T]}\|X(\cdot)\|^{2}_{\mathcal{X}}\right)<\infty.$ For simplicity, we often write $L^{2}_{\mathcal{F}}(0,T;\mathcal{X})$, $H^{1}_{\mathcal{F}}(0,T;\mathcal{X})$ and $L^{\infty}_{\mathcal{F}}(0,T;\mathcal{X})$ whenever no confusion may arise.
Let $T>0$ and $n\in\mathds{N}$, and $G:=(0,1)^{n}\subset\mathbb{R}^{n}$. We consider the following stochastic parabolic equation:
@@
In (1.1), $y$ denotes the state variable, $g$ is a source term representing the intensity of the random force of white noise, and $\gamma^{i},a^{1i}$, and $a^{k}$ (for $i=1,\ldots,n$ and $k=2,3$) are given functions. We refer to Krylov and Rozovskiĭ (1977), Zhou (1992), and Lü and Zhang (2021) for the existence and uniqueness results of (1.1).
The main objective of this paper is to study the inverse random source problem and the Cauchy problem associated with the semi-discrete spatial approximation of the stochastic parabolic equation (1.1). To this end, we introduce the spatial discretization employed throughout this work.
Let $N\in\mathbb{N}$ be sufficiently large and define a set of points $\mathcal{M}=G\cap h\mathbb{Z}^{n}$ with $h=1/(N+1)$, called the primal mesh, we define the dual mesh in the direction $e_{k}$, where $\{e_{k}\}_{k=1}^{n}$ denotes the standard basis of $\mathbb{R}^{n}$, by
$\mathcal{M}_{k}^{\ast}:=\left\{x\pm\frac{h}{2}e_{k}\mid x\in\mathcal{M}\right\}.$
This allows us to define the boundary points in the $e_{k}$ direction as $\partial_{k}\mathcal{M}:=(\mathcal{M}^{\ast}_{k})^{\ast}_{k}\setminus\mathcal{M}$, and we set
$\partial\mathcal{M}:=\bigcup_{k=1}^{n}\partial_{k}\mathcal{M},\quad\textrm{ and }\quad\overline{\mathcal{M}}:=\bigcup_{k=1}^{n}(\mathcal{M}^{\ast}_{k})^{\ast}_{k}.$
@@
(1.2) $\forall x\in\partial_{i}\mathcal{M},\quad\nu_{i}(x):=\begin{cases}1&\mbox{if }x-\tfrac{h}{2}e_{i}\in\mathcal{M}_{i}^{\ast}\ \mbox{and }x+\tfrac{h}{2}e_{i}\notin\mathcal{M}_{i}^{\ast},\\ -1&\mbox{if }x-\tfrac{h}{2}e_{i}\notin\mathcal{M}_{i}^{\ast}\ \mbox{and }x+\tfrac{h}{2}e_{i}\in\mathcal{M}_{i}^{\ast},\\ 0&\mbox{otherwise}.\end{cases}$
@@
(1.3) $\forall x\in\partial_{i}\mathcal{M},\quad t_{r}^{i}(u)(x):=\begin{cases}u\!\left(x-\tfrac{h}{2}e_{i}\right)&\nu_{i}(x)=1,\\ u\!\left(x+\tfrac{h}{2}e_{i}\right)&\nu_{i}(x)=-1,\\ 0&\nu_{i}(x)=0.\end{cases}$
@@
$(D_{k}u)(x):=\frac{\tau_{+k}u(x)-\tau_{-k}u(x)}{h},\quad(A_{k}u)(x):=\frac{\tau_{+k}u(x)+\tau_{-k}u(x)}{2},\quad\text{for }x\in\mathcal{M}^{\ast}_{k},$
where $\tau_{\pm k}u(x):=u\!\left(x\pm\frac{h}{2}e_{k}\right)$.
@@
$\|u\|_{L_{h}^{p}(\mathcal{M})}^{p}:=\int_{\mathcal{M}}|u|^{p}\quad\text{with}\quad\int_{\mathcal{M}}|u|^{p}:=h^{n}\sum_{x\in\mathcal{M}}|u(x)|^{p}.$
Analogously, we denote by $W^{1,p}_{h}(\mathcal{M})$, $W_{h}^{2,p}(\mathcal{M})$, and $L^{\infty}_{h}(\mathcal{M})$ the set of functions defined in $\mathcal{M}$ with the associated norms:
$\|u\|_{W_{h}^{1,p}(\mathcal{M})}^{p}:=\|u\|_{L_{h}^{p}(\mathcal{M})}^{p}+{\sum_{i=1}^{n}}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}u|^{p},$
$\|u\|_{W_{h}^{2,p}(\mathcal{M})}^{p}:=\|u\|_{W_{h}^{1,p}(\mathcal{M})}^{p}+{\sum_{i,j=1}^{n}}\int_{{\mathcal{M}_{ij}^{\ast}}}|D_{ij}^{2}u|^{p},\quad\text{and}\quad\|u\|_{L_{h}^{\infty}(\mathcal{M})}:=\max_{x\in\mathcal{M}}|u(x)|,$
@@
$\int_{\partial\mathcal{M}}u:=\sum_{i=1}^{n}\int_{\partial_{i}\mathcal{M}}u(x)\quad\text{with}\quad\int_{\partial_{i}\mathcal{M}}u(x):=h^{n-1}\sum_{x\in\partial_{i}\mathcal{M}}u(x).$
Then, we denote as $L^{2}_{h}(\partial\mathcal{M})$ and $H^{1}_{h}(\partial\mathcal{M})$ the set of functions defined on $\partial\mathcal{M}$ with the associated norm:
$\left\|u\right\|^{2}_{L^{2}_{h}(\partial\mathcal{M})}:=\int_{\partial\mathcal{M}}|u|^{2}\quad\text{and}\quad\left\|u\right\|^{2}_{H^{1}_{h}(\partial\mathcal{M})}:={\sum_{i=1}^{n}}\left(\int_{\partial_{i}\mathcal{M}}|u|^{2}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\int_{(\partial_{i}\mathcal{M})_{j}^{\ast}}|D_{j}u|^{2}\right),$
respectively. Finally, following Ervedoza and de Gournay (2011), we adopt the discrete $H^{1/2}_{h}$-norm defined for $v\in C(\partial\mathcal{M})$ by
$\|v\|_{H_{h}^{1/2}(\partial\mathcal{M})}:=\min_{\left.u\right|_{\partial\mathcal{M}}=v}\|u\|_{H^{1}_{h}(\mathcal{M})}.$
Following the same reasoning, we analogously define the semi-discrete spaces
$L^{2}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\partial\mathcal{M}))$, $L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M}))$, and
$L^{2}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))$.
Given the preceding definitions, we are now ready to formulate the problems to be addressed. Let $\gamma_{i}$ denote the sampling of $\gamma^{i}$ on both the primal and dual meshes. Furthermore, let $a_{1i}$, $a_{2}$, and $a_{3}$ represent the evaluations of $a^{1i}$, $a^{2}$, and $a^{3}$ on the primal mesh, for $i=1,\ldots,n$, respectively. We now formulate the semi-discrete counterpart of the inverse problems considered in this paper, namely, the semi-discrete inverse random source problem and the semi-discrete Cauchy problem for $n$-dimensional stochastic parabolic equation.
Semi-discrete inverse random source problem. We consider the following $n$-dimensional semi-discrete stochastic parabolic equation, derived from (1.1), assuming $a_{3}=0$:
@@
where the second-order operator is defined by $\mathcal{A}_{h}w:=dw-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(w))dt$, and $w_{0}$ is the initial condition.
The following assumption on $\gamma_{i}$ will be needed throughout the paper. Let $\gamma_{i}>0$ for all i=1,,ni=1,...,n and
$\mbox{reg}(\gamma):=\text{ess}\sup_{\begin{subarray}{c}x\in G\\ i=1,\cdots,n\end{subarray}}\left(\gamma_{i}+\frac{1}{\gamma_{i}}+{\sum_{j=1}^{n}|\partial_{j}\gamma_{i}|^{2}}\right)<c_{0}.$
for any constant $c_{0}>0$. Now, we define the observation operator $\Lambda_{1}:L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))\longrightarrow X_{1}$ as
$\Lambda_{1}(g):=\left(\left.w\right|_{(G_{0}\cap\mathcal{M})\times(0,T)},\left.w\right|_{t=T}\right),$
where $G_{0}$ is an arbitrary open subset of $G$, and $w$ denotes the solution to (1.4) associated with the source term $g$. We set $X_{1}:=L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(G_{0}\cap\mathcal{M}))\times L^{2}_{\mathcal{F}_{T}}(\Omega;L^{2}_{h}(\mathcal{M}))$ endowed with the norm
$\|(f_{1},f_{2})\|_{X_{1}}:=\|f_{1}\|_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(G_{0}\cap\mathcal{M}))}+\|f_{2}\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;L^{2}_{h}(\mathcal{M}))}.$
The direct problem consists in solving (1.4) for a given $g$ and then computing $\Lambda_{1}(g)$. The inverse problem is to recover $g$ from the observation $\Lambda_{1}(g)$. We now state our first main result, which provides a conditional stability estimate for this inverse problem.
@@
Let $a_{1i}\in L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$ for $i=1,\ldots,n$, and $a_{2}\in L_{\mathcal{F}}^{\infty}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))$ where $n^{\ast}$ satisfies
@@
Assume that $g_{k}\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))$ for $k=1,2$, and that, for $i=1,\ldots,n$,
(1.6) $|D_{i}(g_{1}-g_{2})|\leq\mathcal{C}|A_{i}(g_{1}-g_{2})|,\text{ in }{Q_{i}^{*}:=(0,T)\times\mathcal{M}_{i}^{\ast}},\;\mathbb{P}\text{-a.s.}$
Then, there exists a constant $\mathcal{C}>0$, independent of $h$, such that
(1.7) $\|g_{1}-g_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\mathcal{M}))}\leq\mathcal{C}\|\Lambda_{1}(g_{1})-\Lambda_{1}(g_{2})\|_{X_{1}}.$